Give the equation of a conic which is symmetric to the line x+2 = 0.
1
Expert's answer
2020-05-18T19:56:49-0400

The equations of conic sections is simmetric to the line "x+2=0" is

a) Circle "(x+2)^2+(y-k)^2=r^2" center is "(-2,k)," radius is "r"

b) Ellipse "\frac{(x+2)^2}{a^2}+\frac{(y-k)^2}{b^2}=1" center is "(-2,k)" length of major axis is "2a" , length of minor axis is "2b"

c) Hyperbola "\frac{(x+2)^2}{a^2}-\frac{(y-k)^2}{b^2}=1" center is "(-2,k)" distance between the vertices is "2a"

d) Parabola "(x+2)^2=4p(y-k)" vertex is "(-2,k)" , focus is "(-2,k+p)" ,directrix is the line

"y=k-p" , axis is the line "x=-2"